OpenStudy (anonymous):

can some1 tell me if im right There is a positive correlation between the number of times the Striped Ground Cricket chirps per second and the temperature in degrees Fahrenheit. If a scatter plot is made with the number of chirps on the horizontal axis and the trend line is found to be y = 3x + 25, then what would you predict the number of chirps per second to be when the temperature is 55 degrees Fahrenheit? 10 ***** 43 190 I think its the one with the ***

OpenStudy (agent0smith):

is y the temperature...? The question doesn't even specify it :/ it says x is chirps per second, so i guess y is temperature. y = 3x + 25 so at 55 degrees means y=55 55 = 3x+25

OpenStudy (anonymous):

I got 190 when I did it how my teacher told me how to do it what did you get

OpenStudy (agent0smith):

Well the question is written pretty poorly... I'm assuming x is the number of chirps per second based on " the number of chirps on the horizontal axis ", which would make y the temperature. If you solve this: 55 = 3x+25 30 = 3x which gives x=10. If x is actually temperature, then y = 3(55)+25 = 190 But the problem is the question needs to be clearer.

OpenStudy (anonymous):

I agree